If 15.0 g of Ice at -25.0°C is mixed with a mass of antifreeze initially at -3.00°C. If the final temperature of the mixture is -10.0°C, how much antifreeze is present (Cantifreeze=2.24 x 102 J/KgoC)?
1
Expert's answer
2020-12-08T06:33:42-0500

"m_1 = 15.0g = \frac{15}{1000}kg\\
m_2 = ?\\
c_1 = 4.2×10³J/kg.K\\
c_2 = 2.24×10²J/kg.K\\
T_1 = -25.0°C\\
T_2 = -3.0°C\\
T_3 = -10.0°C"


from,

Heat lost = Heat gained

"m_1c_1∆T = m_2c_2∆T\\
m_1c_1(T3-T1) = m_2c_2(T2-T3)"


"\frac{15}{1000}×4.2×10³×(-10+25) = m_2×2.24×10²×(-3+10)\\
63×1000×\frac{15}{1000} = m2×15.68×100"


"m_2 = \dfrac{63×1000×15}{15.68×100×1000}"


"m2 = 0.603kg = 603g"